Summary statistics for returns on two stocks X and Y are listed below.

Mean Variance
Stock X 2.85% 0.005000
Stock Y 5.91% 0.006000

The covariance of returns on stocks X and Y is 0.002800. Consider a portfolio of 30% stock X and 70% stock Y.

What is the mean of portfolio returns?

Please specify your answer in decimal terms and round your answer to the nearest thousandth (e.g., enter 12.3 percent as 0.123).
Answer #1

for return of the portfolio R=0.3X+0.7Y

therefore  mean of portfolio returns E(R) =0.3E(X)+0.7*E(Y)=0.3*2.85+0.7*5.91 =4.992 % ~ 0.050